OpenStudy (anonymous):

How do you foil with this problem: x (3x-4x squared)

OpenStudy (anonymous):

\[ x (3x-4x^2)\]is this it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you can foil out another x because there is an x in 3x and in 4x^2

OpenStudy (anonymous):

How do i foil?

OpenStudy (anonymous):

sorry i don't remember from algebra... trying to study for my praxis

OpenStudy (anonymous):

you are basically pulling out the x from each of the x and you are pulling it out of the parentheses \[x(3x−4x^2)\]\[x(3x−4x*x)\]\[x*x(3−4x)\]

OpenStudy (anonymous):

see how 4x^2 had two x and even if we pulled out one x there is still one left inside while 3x only had one x so when we pulled it outside it went away

OpenStudy (anonymous):

okay

OpenStudy (phi):

FOIL is used to remember how to multiply two binomials (x+y)(a+b) It is short for first (multiply the first terms together xa Outer x*b Inner y*a Last y*b then add together xa+xb+ya+by for your problem, you do not have two binomials You can use the distributive law: x(a+b)= x*a+x*b in your problem x (3x-4x^2) = 3x^2-4x^3